OpenStudy (anonymous):

What is the length of the hypotenuse of the right triangle? a=16 b=4 c=c Help please!

OpenStudy (anonymous):

@Mr.Man420 @sleepyjess @flasht

OpenStudy (anonymous):

A^2+B^2=C^2 256+16=c^2

OpenStudy (anonymous):

These are the choices A. 400 B. 8.9 C.16.5 D. 4.5

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Are you positive? Lol

OpenStudy (anonymous):

256+16=272 and the square root of 272 is 16.5 so im pretty sure(:

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

No problem what so ever thank you(: